Skip to main content

Problem 107: In a given city, 6% of all drivers get at least one parking ticket per year. Use the Poisson approximation to the binomial distribution to determine the probabilities that among 80 drivers randomly chosen in this given city:
a. 4 will get at least one parking ticket in any given year,
b. At least 3 will get at least at least one parking ticket in any given year,
c. Anywhere from 3 to 6, inclusive will get at least one parking ticket in any given year.

Problem 107: In a given city, 6% of all drivers get at least one parking ticket per year. Use the Poisson approximation to the binomial distribution to determine the probabilities that among 80 drivers randomly chosen in this given city:
a. 4 will get at least one parking ticket in any given year,
b. At least 3 will get at least at least one parking ticket in any given year,
c. Anywhere from 3 to 6, inclusive will get at least one parking ticket in any given year.


Solution:


Poisson distribution, $f(x;\lambda )=\frac{{\lambda }^{x}{e}^{-\lambda }}{x!}$

a. 4 will get at least one parking ticket in any given year
Given that $x=4, n=80, p=0.06, np = 4.8$
$f(4;4.8)=\frac{{4.8}^{4}{e}^{-4.8}}{4!} = 0.182029$


b. At least 3 will get at least at least one parking ticket in any given year
Given that $x\ge 3, n=80, p=0.06, np = 4.8$
$P(x\ge 3) = 1-P(x<3) = 1-[P(x=0)+P(x=1)+P(x=2)]$
$P(x=0) = f(0;4.8)=\frac{{4.8}^{0}{e}^{-4.8 }}{0!} = 0.00823$
$P(x=1) = f(1;4.8)=\frac{{4.8}^{1}{e}^{-4.8 }}{1!} = 0.039503$
$P(x=2) = f(2;4.8)=\frac{{4.8}^{2}{e}^{-4.8 }}{2!} = 0.094807$
$P(x\ge 3) = 1-[P(x=0)+P(x=1)+P(x=2)] = 1-(0.00823+0.039503+0.094807) = 1-0.1425 =0.857.$

c. Anywhere from 3 to 6, inclusive will get at least one parking ticket in any given year
Given that $3 \le x\le 6, n=80, p=0.06, np = 4.8$
$P(3 \le x\le 6) = P(x=3)+P(x=4)+P(x=5)+P(x=6)]$
$P(x=3 = f(3;4.8)=\frac{{4.8}^{0}{e}^{-4.8 }}{0!} = 0.151691$
$P(x=4) = f(4;4.8)=\frac{{4.8}^{1}{e}^{-4.8 }}{1!} = 0.182029$
$P(x=5) = f(5;4.8)=\frac{{4.8}^{3}{e}^{-4.8 }}{3!} = 0.174748$
$P(x=6) = f(6;4.8)=\frac{{4.8}^{4}{e}^{-4.8 }}{4!} = 0.139798$
$P(3 \le x\le 6) = P(x=3)+P(x=4)+P(x=5)+P(x=6)] = 0.162$

Comments

Popular posts from this blog

Problem 105: If a bank receives on the average α = 6 bad checks per day, what are the probabilities that it will receive (a) 4 bad checks on any given day?
(b) 10 bad checks over any 2 consecutive days?

Problem 105: If a bank receives on the average α = 6 bad checks per day, what are the probabilities that it will receive (a) 4 bad checks on any given day? (b) 10 bad checks over any 2 consecutive days? Solution: Poisson distribution, $f(x;\lambda )=\frac{{\lambda }^{x}{e}^{-\lambda }}{x!}$ a. 4 bad checks on any given day Given that $x=4, \lambda = \alphaT = 6\dot 1 = 6$ $f(4;6 )=\frac{{6 }^{4}{e}^{-6 }}{4!} = 0.134$ b. 10 bad checks over any 2 consecutive days Given that $x=10, \lambda = \alphaT = 6\dot 2 = 12$ $f(10;12 )=\frac{{12 }^{10}{e}^{-12 }}{10!} = 0.104837$

Problem 101: Among the 300 employees of a company, 240 are union members, while the others are not, If 8 of the employees are chosen to serve on the administrative committee, find the probability that 5 of them will be union member while the others are not.
a. the formula for the hypergeometric distribution;
b. the formula for the binomial distribution as an approximation.

Problem 101: Among the 300 employees of a company, 240 are union members, while the others are not, If 8 of the employees are chosen to serve on the administrative committee, find the probability that 5 of them will be union member while the others are not. a. the formula for the hypergeometric distribution; b. the formula for the binomial distribution as an approximation. Solution: a. the formula for the hypergeometric distribution Given that $x=5,n=8,a=240, N=300$ Hypergeometric distribution, $h(x;n,a,N)=\frac{\left(\begin{array}{c}a\\ x\end{array}\right)\left(\begin{array}{c}N-a\\ n-x\end{array}\right)}{\left(\begin{array}{c}N\\ n\end{array}\right)}$ $P(x=5) = h(5;8,240,300)=\frac{\left(\begin{array}{c}240\\ 5\end{array}\right)\left(\begin{array}{c}300-240\\ 8-5\end{array}\right)}{\left(\begin{array}{c}300\\ 8\end{array}\right)} = 0.1470.$ b. the formula for the binomial distribution as an approximation? Given that $x=5,n=8,p = \frac{240}{300}$ $P(x= 5) =...

Problem 91: Suppose that a lot of 5000 electrical fuses contains 5% defectives. If a sample of 5 fuses is tested, find the probability of observing at least one defective.

Problem 91: Suppose that a lot of 5000 electrical fuses contains 5% defectives. If a sample of 5 fuses is tested, find the probability of observing at least one defective. Solution: Given that $n=5,$ and $p=0.05$ at least one defective; $P(x\ge 1) = 1-P(x=0)$ $b(x;n,p)=\left(\begin{array}{c}n\\ x\end{array}\right){p}^{x}(1-p)^{n-x}$ $P(x\ge 1) = 1 - b(0;5,0.05) = 1 - \left(\begin{array}{c}5\\ 0\end{array}\right){0.05}^{0}(1-0.05)^{5-0} = 1-(0.95)^5 = 0.226.$